Please explain with example: follow the comment:

If A is an n×n matrix with the property that Ax = 0
for all x ∈ Rn, show that A = O. Hint: Let x = ej
for j = 1, . . . , n.Ax = 0 for all XEO Let A-(a,a,. Let e.-| | | ← jth element Comments (1) Anonymous why there is 1 in side ej Submit Step 2 of 2 Hence Ae,-a Since AX-0 for all Xe0 Ae, = 0 (Since e, e。) a0 for 1sjS

Why ej=(1,....n) then it comes out it is a column vector and all zero except 1 inside, i don't get it
Answer #1

For n = 2 then A-(0111 012 a21 a22 x E R2 means x = where 1, r2 ER. where R-( 0 and ER2 0

0 Choose x- en 0 and 0 0 From (1) (a «i2 (1 a11 01 a21 0 (3) a11 a21 0 0

0 0 a12 01 C21 22 0 122 艹a12 = 0, a22 = 0 (4)

Hence from (3) and (4), A -0 0y ei and We denote the first colunrn of A i.e. So from (3) and (4, a 0, a2-0 0 We generally den

For n-3, then T1 T3 an a12 a13 A -| a21 a22 a23 Z31 132 33/ 0 Ax-0- xER3 (given) 0 Then take xe e and e3 respectively as n 2 where

11 12 13 a22 a31 32 33 Nou, LetA-- | a21 a23 (al , a), as) a12 13 where, a a31 132 133

Since Ax 0 x ER so

Again

0 131 a12 0 132 a13 133

a1, a2, ..., aj, ...., an a1j (42 Cl where, aj = li 21 an then Aej = 0, j = 1(1)n where

YO 0 0 0 0 0 0 , e2 0 0 0 Note that 1 appears in e, at jth position. then aj -0Vj1)n Hence A=0

Alternative proof:

Let A is a matrix of order n×n (ie, A has n rows and n columns) then we write A as Z11 l12 a1j aln 21 122 2j d2n Zil 4i2 aij... ain an1 an2 Inj ann/ Given, 21 n×1 where, X 7t column vector which has n components (since x E R), 0 0 0

Let A is a matrix of order n×n (ie, A has n rows and n columns) then we write A as 0 Ax= =0= 0 Since Ax-0 x E Rn 0 Now we choose x1 = 1, x2-x3 = . . . = xn = 0 then E R 0 0 or, ail = 0V-1(1)a i.e. all = a21-. . . _ an 1 = 0 (2)

0 Ņow we choose x1 = 0,T2=1, =x3 = . . . = Zn = 0 then E R 0 0 or, a0Vi 1(1)n i.e. a12 a22-... -an20.... (3) Proceeding in this way, we can show that for the choice r.rit1 _ ...-.rn- and lastly for the choice Combining (2) - (5), we get A =0 = matrix of order n × n whose all elements are zero